# DenoiseReID: Denoising Model for Representation Learning of Person Re-Identification

### v1

@@ 4 Experiments
4.1 FEFDFA Performance Analysis
Table 1: FEFDFA performance analysis using mAP metric. The baseline method TransReID-SSL is based on ViT-small backbone. "label-arg" means label-argumented and "merged ds" means merged datasets.
@@ 4.1 FEFDFA Performance Analysis
+FEFDFA (lable-free) 80.92 (↑ 0.52) 66.87 (↑ 0.57) 91.82 (↑ 0.62) 83.72 (↑ 0.22)
+FEFDFA (label-arg) 81.22 (↑ 0.82) 67.33 (↑ 1.03) 92.05 (↑ 0.85) 84.11 (↑ 0.61)
+FEFDFA (merged ds) 80.98 (↑ 0.58) 66.99 (↑ 0.69) 91.80 (↑ 0.60) 83.86 (↑ 0.36)
As mentioned in Section 3.3, Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) is an unsupervised denoising module, and its training does not require the assistance of label information. We conducte the following experiments to identify three key issues.
(1) Is this label-free and unsupervised training denoising plugin effective? As shown in Table 1 (line2), compared with baseline method (line1), the baseline method performs better after adding our label-free plugin, which shows that our FEFDFA does have denoising capability for features.
(2) Could introducing label information for supervised training further improve performance? Introducing label information is actually adding $Loss_{ReID}$ as mentioned in Section 3.4 as a supervised signal. As shown in Table 1 (line3), baseline method with label-argumented FEFDFA achieve improvements of 0.32% - 0.70% on the mAP metric, indicating that our denoising plugin has label compatibility, in other words, the plug-in is effective for feature denoising regardless of label-argumented supervised or lable-free unsupervised training.
(3) Since our plugin can perform unsupervised denoising of features, it is natural to think about whether adding more data for training the plugin could further improve its performance? We merge four datasets for training and then test on each dataset using mAP to evaluate. Comparing the results of training on sigle dataset (line2) with on merged datasets (line4), we found that adopting other datasets for unsupervised learning can further improve the performance of FEFDFA, which also proves that FEFDFA has good generalization ability.
To demonstrate that our method can perform unsupervised learning and has good generalization, we merged four datasets and rearranged the sequence IDs to ensure the reliability of the experiment. The model is tested on the entire dataset. During the training process, we freeze the baseline parameters and only train the DenoiseReID module, without the need for labels, for unsupervised learning. Then test on a single dataset and compare the results of training on a single dataset. As shown in Table 1, it can be observed that adding unlabeled training data from different datasets can improve the model’s performance on a single dataset, proving that this module has a certain degree of generalization.
@@ 4.2 Comparison with State-of-the-Art ReID Methods
We compare several state-of-the-art ReID methods on four datasets. One of the best performing comparison methods is TransReID-SSL, which is a series of ReID methods based on the ViT backbones. Other methods are based on structures such as CNNs. We add the proposed FEFDFA to TransReID-SSL series and observe their performance. As shown in Table 2, we have the following findings:
Table 2: Comparison with State-of-the-Art ReID methods using mAP and Rank-1 metrics.
@@
TransReID+FEFDFA ViT-base-ics 68.10 85.72 89.56 95.50 82.35 90.87 84.15 86.39
TransReID+FEFDFA ViT-base 62.23 82.02 87.25 94.63 80.12 89.33 82.44 84.61
TransReID-SSL+FEFDFA ViT-small 67.33 85.50 92.05 96.68 81.22 88.72 84.11 86.47
TransReID-SSL+FEFDFA ViT-base 75.35 89.62 93.26 96.55 84.31 92.90 88.08 89.29
@@
(2) The methods using our plugin outperforms the original methods with the same backbone on all datasets. In addition, the performance improvement of small-scale backbones with the addition of FEFDFA is more significant than the large-scale backbones approach due to the fact that FEFDFA is essentially a denoising module that removes the noise contained in the features during the inference stage. For large-scale backbones, the extracted features already have good performance, so the denoising amplitude is limited. It has already fitted the dataset well. For small-scale backbones with poor performance, due to their limited fitting ability, there is a certain amount of noise in the extracted features during the inference stage. Denoising them can obtain better feedback.
(3) In fact, our FEFDFA can be applied to any other backbone, just add it to each layer. In particular, the performance improvement of adding the denoising plugin to a poorly performing backbone might be more significant. This needs to be further verified in subsequent work. However, it is undeniable that we have verified the denoising ability of the FEFDFA in the currently optimal ReID method.
@@
4.3 Parameter Fusion Performance Analysis
Our Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) is a computation-free way, In section 3.3, we proved by theoretical derivation that inserting our plugin into each layer and fusion it does not introduce additional computation. In this section, we also conduct related validation experiments, the results of which are shown in Table 3.
Table 3: Parameter Fusion Performance Analysis. The Feature Extraction and Feature Denoising Unified Framework (FEFDUF) denoises based on the features of the final layer, while the Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) denoises based on the features of each layer. The baseline method TransReID-SSL is based on ViT-small backbone.
@@ 4.3 Parameter Fusion Performance Analysis
+FEFDUF 80.76% 66.81% 91.07% 83.59% 0.39s (+15%)
+FEFDFA 81.22% 67.33% 92.05% 84.11% 0.34s (+0%)
Compare to the baseline method TransReID-SSL, adding FEFDUF is able to improve the the performance, proving that feature based denoising is effective. However, it also brings extra inference latency (about 15%) because it is adding an extra parameter-independent denoising module at the end of the model.
Adopting FEFDFA achieves a greater increase, it denoise the features on each layer, which can better remove noise at each stage because the noise in the inference stage comes from multiple aspects, which may be the noise in the input image or generated when passing through the network. Denoising each layer avoids noise accumulation and obtains a better quality output. Most importantly, since the operation of fusion can merge the parameters of the denoising module with the original parameters, the adoption of FEFDFA does not take extra inference latency cost, which is a computation-free efficient approach.
@@ 4.4 Experiments on Classification Tasks
Method
Datasets
param
@@
acc@1(ours)
acc@5(ours)
SwinV2-T [36] ImageNet-1k 28M 81.82% 95.88% 82.13% 96.06%
Vmanba-T [35] ImageNet-1k 30M 82.38% 95.80% 82.51% 95.89%
ResNet50 [18] ImageNet-1k 26M 76.13% 92.86% 76.28% 92.95%
@@
The Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA) is based on denoising at the feature level and has good generalization. To demonstrate its good generalization ability, we conduct experiments on other vision tasks to test the effectiveness of the FEFDFA. We validated the generalization ability of FEFDFA in image classification tasks on ImageNet-1k [11] datasets and three fine-grained image classification datasets (CUB200 [50], Oxford-Pet [42], and Flowers [41]). The accuracy index is selected as the evaluation indicator to evaluate the performance of the model. As shown in Table 4, We compared multiple classic backbones for representation learning on ImageNet-1k, and after adding the FEFDFA module, the accuracy of both top-1 and top-5 metrics improved without adding model parameters. As shown in Table 4, our method shows significant improvement in accuracy metrics compared to baseline on three fine-grained classification datasets. Prove that the FEFDFA module can improve the model’s ability in image classification for different classification tasks.It has been proven that our method can enhance the model’s representation learning ability and obtain more effective features through denoising without adding additional time cost. More experimental analysis can be found in Table 6 in Section C of the appendix.
@@ 5 Conclusion
In this work, we demonstrated that the diffusion model paradigm is effective for feature level denoising in discriminative model, and proposed computation-free and label-free algorithm: Feature Extraction and Feature Denoising Fusion Algorithm (FEFDFA). It utilizes the denoising ability of diffusion models to denoise the features in the feature extraction layer, and fuses the parameters of the denoising layer with those of the feature extraction layer through parameter fusion, further improving retrieval accuracy without incurring additional computational costs. We validated the effectiveness of the FEFDFA method on multiple common image discrimination task datasets.
@@ Appendix A Experimental settings
Datasets and Evaluation metrics. We conducted training and evaluation on four datasets: DukeMTMC-reID [65], Market-1501 [63], MSMT17 [56], and CUHK-03 [30]. These datasets encompass a wide range of scenarios for person re-identification. For accuracy, we use standard metrics including Rank-1 curves (The probability that the image with the highest confidence in the search results is the correct result.) and mean average precision (MAP). All the results are from a single query setting.
Implementation Details. We implemented our method using Python on a server equipped with a 2.10GHz Intel Core Xeon (R) Gold 5218R processor and two NVIDIA RTX 3090 GPUs. The epochs we trained are set to 120, the learning rate is set to 0.0004, the batch size during training is 64, the inference stage is 256, and the diffusion step size $t$ is set to 1000.
Training and evaluation. To better constrain the performance of the denoised features of the FEFDFA for downstream tasks, we used alternate fine-tune methods. The parameters of the FEFDFA and baseline were trained alternately, and when training a part of the parameters, the rest of the parameters were frozen and fine-tuned for 10 epochs at a time, with a total number of epochs of 120. When evaluating, we averaged the results of the experiments under the same settings for 5 times, thus ensuring the reliability of the data.
@@ Appendix B Experiment on Vehicle Identification
In the image retrieval task, we also conducted experiments to verify the effectiveness of our method on the vehicle recognition task. Vehicle recognition in practical scenarios often results in images containing a large amount of noise due to environmental factors such as lighting or occlusion, which increases the difficulty of detection. Our method is based on denoising to obtain features with better representation ability. Therefore, we want to experimentally verify whether the FEFDFA module plays a role in vehicle recognition tasks with higher noise levels. We selected vehicleID [34] as the dataset, vehicle-ReID [10] as the baseline, and ResNet50 as the feature extractor for the experiment.
Table 5: The performance of the FEFDFA module in vehicle recognition tasks.
@@
vehicle-ReID ResNet50 vehicleID 76.4% 69.1%
vehicle-ReID(FEFDFA) ResNet50 vehicleID 77.3% 70.2%
From the results in Table 5, it can be seen that FEFDFA has excellent performance in vehicle detection tasks. Compared to the baseline, adding the FEFDFA module significantly improves both mAP and Rank-1 metrics without incurring additional detection time costs. Verified the denoising ability of the FEFDFA module in noisy environments.
@@ Appendix C Experiment on Large Scale Image Classification Tasks
In this section, we aim to test the generalization ability of FEFDFA in other tasks. We conducted experiments on two image classification datasets, ImageNet-1k and Cifar-10. These two datasets are both classic image classification datasets, rich in common images in daily life, and belong to large-scale image databases. ImageNet-1k is a subset of the ImageNet dataset, containing images from 1000 categories. Each category typically has hundreds to thousands of images, totaling over one million images. The Cfiar-10 contains 60000 32x32 pixel color images, divided into 10 categories. Each category contains 6000 images. To evaluate the effectiveness of our method, we use standard metrics, including Top-1 accuracy and Top-5 accuracy, which are commonly used to evaluate model performance in image classification tasks and are widely used on various image datasets, and we conducted detailed comparative experiments on multiple backbones and models with different parameter versions to verify the reliability of our method.
@@
Method
Datasets
param
@@
acc@1(ours)
acc@5(ours)
SwinV2-T [36] ImageNet-1k 28M 81.82% 95.88% 82.13% 96.06%
SwinV2-S [36] ImageNet-1k 50M 83.73% 96.62% 83.97% 96.86%
SwinV2-B [36] ImageNet-1k 88M 84.20% 96.93% 84.31% 97.06%
Vmanba-T [35] ImageNet-1k 30M 82.38% 95.80% 82.51% 95.89%
Vmanba-S [35] ImageNet-1k 50M 83.12% 96.04% 83.27% 96.22%
Vmanba-B [35] ImageNet-1k 89M 83.83% 96.55% 83.91% 96.70%
ResNet50 [18] ImageNet-1k 26M 76.13% 92.86% 76.28% 92.95%
@@
As shown in Table 6, We compared multiple classic backbones for representation learning on two datasets, and after adding the FEFDFA module, the accuracy metrics improved without adding model parameters. It has been proven that our method can enhance the model’s representation learning ability and obtain more effective features through denoising without increasing additional time costs. Moreover, this method can be well generalized to other image tasks.
Appendix D Limitations

### v2

@@ 4 Experiments
{adjustwidth}
-1cm-1cm
Table 1: Experimental results on various discriminative tasks.
Task
Model
Backbone
Dataset
Metric
Baseline
+DenoiseRep
Classification SwinT [36] SwinV2-T ImageNet-1k acc@1 81.82% 82.13%
Person-ReID TransReID-SSL [37] ViT-S MSMT17 mAP 66.30% 67.33%
Detection Mask-RCNN [he2017mask] SwinV2-T COCO AP 42.80% 44.30%
Segmentation FCN [long2015fully] ResNet-50 ADE20K BIoU 28.70% 29.90%
Our proposed DenoiseRep is a versatile method that can be incrementally applied to various discriminative tasks. Table 1 demonstrates that DenoiseRep yields stable and substantial improvements across image classification, object detection, image segmentation, and person re-identification. Given that person re-identification is a nuanced image retrieval task that poses a greater challenge to feature discriminability, we take it as our benchmark for model analysis. Details of the experimental settings are provided in Appendix A. Additional experimental results on various tasks are presented in Appendices B, C, D, and E.
4.1 Analysis of Label Informations
Table 2: DenoiseRep is a label-free method that can also be effectively complemented with labels when they are available. The table below analyzes the effectiveness of using labels. The baseline method, TransReID-SSL, is based on a ViT-small backbone. "Label-free" indicates training without labels, "label-augmented" refers to the use of labels, and "merged dataset" denotes the use of combined datasets without labels.
@@ 4.1 Analysis of Label Informations
+DenoiseRep (label-free) 80.92 (↑ 0.52) 66.87 (↑ 0.57) 91.82 (↑ 0.62) 83.72 (↑ 0.22)
+DenoiseRep (label-aug) 81.22 (↑ 0.82) 67.33 (↑ 1.03) 92.05 (↑ 0.85) 84.11 (↑ 0.61)
+DenoiseRep (merged ds) 80.98 (↑ 0.58) 66.99 (↑ 0.69) 91.80 (↑ 0.60) 83.86 (↑ 0.36)
As mentioned in Section 3.3, DenoiseRep is an unsupervised denoising module, and its training does not require the assistance of label information. We conducte the following experiments to identify three key issues.
(1) Is this label-free and unsupervised training denoising plugin effective? As shown in Table 2 (line2), compared with baseline method (line1), the baseline method performs better after adding our label-free plugin, which shows that our method does have denoising capability for features.
(2) Could introducing label information for supervised training further improve performance? Introducing label information is actually adding $Loss_{l}$ as mentioned in Section 3.4 as a supervised signal. As shown in Table 2 (line3), baseline method with label-argumented DenoiseRep achieve improvements of 0.32% - 0.70% on the mAP metric, indicating that our denoising plugin has label compatibility, in other words, the plug-in is effective for feature denoising regardless of label-argumented supervised or lable-free unsupervised training.
(3) Since our plugin can perform unsupervised denoising of features, it is natural to think about whether adding more data for training the plugin could further improve its performance? We merge four datasets for training and then test on each dataset using mAP to evaluate. Comparing the results of training on sigle dataset (line2) with on merged datasets (line4), we found that adopting other datasets for unsupervised learning can further improve the performance of DenoiseRep, which also proves that DenoiseRep has good generalization ability.
To demonstrate that our method can perform unsupervised learning and has good generalization, we merged four datasets and rearranged the sequence IDs to ensure the reliability of the experiment. The model is tested on the entire dataset. During the training process, we freeze the baseline parameters and only train the DenoiseRep module, without the need for labels, for unsupervised learning. Then test on a single dataset and compare the results of training on a single dataset. As shown in Table 2, it can be observed that adding unlabeled training data from different datasets can improve the model’s performance on a single dataset, proving that this module has a certain degree of generalization.
@@ 4.2 Comparison with State-of-the-Art ReID Methods
We compare several state-of-the-art ReID methods on four datasets. One of the best performing comparison methods is TransReID-SSL, which is a series of ReID methods based on the ViT backbones. Other methods are based on structures such as CNNs. We add our method to TransReID-SSL series and observe their performance. As shown in Table 3, we have the following findings:
Table 3: Comparison with state-of-the-art ReID methods.
@@
CLIP-REID [li2023clip] ViT-base 75.80 89.70 90.50 95.40 83.10 90.80
TransReID + DenoiseRep ViT-base-ics 68.10 85.72 89.56 95.50 82.35 90.87 84.15 86.39
TransReID + DenoiseRep ViT-base 62.23 82.02 87.25 94.63 80.12 89.33 82.44 84.61
TransReID-SSL + DenoiseRep ViT-small 67.33 85.50 92.05 96.68 81.22 88.72 84.11 86.47
TransReID-SSL + DenoiseRep ViT-base 75.35 89.62 93.26 96.55 84.31 92.90 88.08 89.29
CLIP-REID + DenoiseRep ViT-base 76.30 90.60 91.10 95.80 83.70 91.60
@@
(2) The methods using our plugin outperforms the original methods with the same backbone on all datasets. In addition, the performance improvement of small-scale backbones with the addition of DenoiseRep is more significant than the large-scale backbones approach due to the fact that DenoiseRep is essentially a denoising module that removes the noise contained in the features during the inference stage. For large-scale backbones, the extracted features already have good performance, so the denoising amplitude is limited. It has already fitted the dataset well. For small-scale backbones with poor performance, due to their limited fitting ability, there is a certain amount of noise in the extracted features during the inference stage. Denoising them can obtain better feedback.
(3) In fact, our method can be applied to any other backbone, just add it to each layer. In particular, the performance improvement of adding the denoising plugin to a poorly performing backbone might be more significant. This needs to be further verified in subsequent work. However, it is undeniable that we have verified the denoising ability of the DenoiseRep in the currently optimal ReID method.
@@
4.3 Analysis of Parameter Fusion
The proposed DenoiseRep is computation-free. In section 3.3, we proved by theoretical derivation that inserting our denoising layer into each feature layer and fusion it does not introduce additional computation. In this section, we also conduct related validation experiments, the results of which are shown in Table 4.
Table 4: Parameter Fusion Performance Analysis. The $\textit{DenoiseRep}^{-}$ denoises based on the features of the final layer, while the DenoiseRep denoises based on the features of each layer. The baseline method TransReID-SSL is based on ViT-small backbone.
@@ 4.3 Analysis of Parameter Fusion
+$\textit{DenoiseRep}^{-}$ 80.76% 66.81% 91.07% 83.59% 0.39s (+15%)
+DenoiseRep 81.22% 67.33% 92.05% 84.11% 0.34s (+0%)
Compare to the baseline method TransReID-SSL, adding $\textit{DenoiseRep}^{-}$ is able to improve the the performance, proving that feature based denoising is effective. However, it also brings extra inference latency (about 15%) because it is adding an extra parameter-independent denoising module at the end of the model.
Adopting DenoiseRep achieves a greater increase, it denoise the features on each layer, which can better remove noise at each stage because the noise in the inference stage comes from multiple aspects, which may be the noise in the input image or generated when passing through the network. Denoising each layer avoids noise accumulation and obtains a better quality output. Most importantly, since the operation of fusion can merge the parameters of the denoising module with the original parameters, the adoption of DenoiseRep does not take extra inference latency cost, which is a computation-free efficient approach.
@@ 4.4 Experiments on Classification Tasks
Method Datasets Param
@@
Baseline +DenoiseRep Baseline +DenoiseRep
SwinV2-T [36] ImageNet-1k 28M 81.82% 82.13% 95.88% 96.06%
Vmanba-T [35] ImageNet-1k 30M 82.38% 82.51% 95.80% 95.89%
ResNet50 [18] ImageNet-1k 26M 76.13% 76.28% 92.86% 92.95%
@@
The DenoiseRep is based on denoising at the feature level and demonstrates strong generalization capabilities. To validate this generalization ability, we conduct experiments on other vision tasks to test the effectiveness of the DenoiseRep. We validate the generalization ability of DenoiseRep in image classification tasks on ImageNet-1k [11] datasets and three fine-grained image classification datasets (CUB200 [50], Oxford-Pet [42], and Flowers [41]). The accuracy index is chosen as the evaluation metric to assess model performance.
As shown in Table 5, we compare multiple classic backbones for representation learning on ImageNet-1k, and after adding the DenoiseRep, the accuracy of both top-1 and top-5 metrics improve without adding model parameters. Our method shows significant improvement in accuracy metrics compared to baseline on three fine-grained classification datasets. Prove that the DenoiseRep can improve the model’s ability in image classification for different classification tasks. Additionally, our method proves to enhance the model’s representation learning ability and extract more effective features through denoising without incurring additional time costs. More experimental analysis can be found in Table 7 in Section C of the appendix.
@@ 5 Conclusion
In this work, we demonstrate that the diffusion model paradigm is effective for feature level denoising in discriminative model, and propose a computation-free and label-free method: DenoiseRep. It utilizes the denoising ability of diffusion models to denoise the features in the feature extraction layer, and fuses the parameters of the denoising layer and the feature extraction layer, further improving retrieval accuracy without incurring additional computational costs. We validate the effectiveness of the DenoiseRep method on multiple common image discrimination task datasets.
Acknowledgement
This work was supported by the National Natural Science Foundation of China (62202041), National Key Research and Development Program of China under Grant (2023YFC3310700) and Fundamental Research Funds for the Central Universities (2023JBMC057).
@@ Appendix A Experimental settings
Datasets and Evaluation metrics. We conduct training and evaluation on four datasets: DukeMTMC-reID [65], Market-1501 [63], MSMT17 [56], and CUHK-03 [30]. These datasets encompass a wide range of scenarios for person re-identification. For accuracy, we use standard metrics including Rank-1 curves (The probability that the image with the highest confidence in the search results is the correct result.) and mean average precision (MAP). All the results are from a single query setting.
Implementation Details. We implement our method using Python on a server equipped with a 2.10GHz Intel Core Xeon (R) Gold 5218R processor and two NVIDIA RTX 3090 GPUs. The epochs we trained are set to 120, the learning rate is set to 0.0004, the batch size during training is 64, the inference stage is 256, and the diffusion step size $t$ is set to 1000.
Training and evaluation. To better constrain the performance of the denoised features of the DenoiseRep for downstream tasks, we employ alternating fine-tune methods. The parameters of the DenoiseRep and baseline are trained alternately, and when training a part of the parameters, the rest of the parameters are frozen and fine-tuned for 10 epochs at a time, with a total number of epochs of 120. When evaluating, we average the results of the experiments under the same settings for 5 times, thus ensuring the reliability of the data.
@@ Appendix B Experiment on Vehicle Identification
In the image retrieval task, we also conduct experiments to verify the effectiveness of our method on the vehicle recognition task. Vehicle recognition in practical scenarios often results in images containing a large amount of noise due to environmental factors such as lighting or occlusion, which increases the difficulty of detection. Our method is based on denoising to obtain features with better representation ability. Therefore, we want to experimentally verify whether the DenoiseRep plays a role in vehicle recognition tasks with higher noise levels. We select vehicleID [34] as the dataset, vehicle-ReID [10] as the baseline, and ResNet-50 as the feature extractor for the experiment.
Table 6: The performance of the DenoiseRep on vehicle recognition tasks.
@@
vehicle-ReID ResNet-50 vehicleID 76.4% 69.1%
vehicle-ReID + DenoiseRep ResNet-50 vehicleID 77.3% 70.2%
From the results in Table 6, it can be seen that DenoiseRep demonstrates excellent performance in vehicle detection tasks. Compared to the baseline, adding the DenoiseRep significantly improves both mAP and Rank-1 metrics without incurring additional detection time costs. It verifies the denoising ability of the DenoiseRep in noisy environments.
@@ Appendix C Experiment on Large Scale Image Classification Tasks
In this section, we aim to test the generalization ability of DenoiseRep in other tasks. We conduct experiments on two image classification datasets, ImageNet-1k and Cifar-10. These two datasets are both classic image classification datasets, rich in common images in daily life, and belong to large-scale image databases. ImageNet-1k is a subset of the ImageNet dataset, containing images from 1000 categories. Each category typically has hundreds to thousands of images, totaling over one million images. The Cfiar-10 contains 60000 32x32 pixel color images, divided into 10 categories. Each category contains 6000 images. To evaluate the effectiveness of our method, we use standard metrics, including Top-1 accuracy and Top-5 accuracy, which are commonly used to evaluate model performance in image classification tasks and are widely used on various image datasets, and we conduct detailed comparative experiments on multiple backbones and models with different parameter versions to verify the reliability of our method.
@@
Method Datasets Param
@@
Baseline +DenoiseRep Baseline +DenoiseRep
SwinV2-T [36] ImageNet-1k 28M 81.82% 82.13% 95.88% 96.06%
SwinV2-S [36] ImageNet-1k 50M 83.73% 83.97% 96.62% 96.86%
SwinV2-B [36] ImageNet-1k 88M 84.20% 84.31% 96.93% 97.06%
Vmanba-T [35] ImageNet-1k 30M 82.38% 82.51% 95.80% 95.89%
Vmanba-S [35] ImageNet-1k 50M 83.12% 83.27% 96.04% 96.22%
Vmanba-B [35] ImageNet-1k 89M 83.83% 83.91% 96.55% 96.70%
ViT-S [14] ImageNet-1k 22M 83.87% 84.02% 96.73% 96.86%
ViT-B [14] ImageNet-1k 86M 84.53% 84.64% 97.15% 97.23%
ResNet50 [18] ImageNet-1k 26M 76.13% 76.28% 92.86% 92.95%
@@
As shown in Table 7, we compare multiple classic backbones for representation learning on two datasets, and after adding the DenoiseRep, the accuracy metrics improves without adding model parameters. Our method demonstrates the capability to enhance the model’s representation learning ability and extract more effective features through denoising, all while maintaining the same time costs. Moreover, DenoiseRep generalizes effectively to image classification tasks.
Appendix D Experiment on Image Detection Task
In this section, we aim to test the generalization ability of DenoiseRep in image detection tasks. We conduct experiments on the COCO [karpathy2015deep] dataset. The COCO (Common Objects in Context) dataset is a widely used dataset for large-scale image recognition, object detection, and image segmentation, particularly in computer vision tasks. It contains 80 types of objects, such as people, animals, daily necessities, etc., covering various common items in daily life. To verify that our method is model independent, we conduct experiments using different models including Mask-RCNN [he2017mask], Faster-RCNN [44], ATSS [zhang2020bridging], YOLO [redmon2018yolov3], DETR [carion2020end] and CenterNet [zhou2019objects], as well as diverse backbones. To evaluate the effectiveness of our method, we used standard metrics including AP, AP50, and AP75, which are commonly used metrics in object detection tasks to evaluate model performance, particularly for evaluating the effectiveness of bounding box detection. It is also an important part of the COCO dataset evaluation criteria, which can measure the detection ability of the model in multiple categories and scales.
Table 8: The effectiveness of our method in image detection tasks was validated on COCO.
Methods Backbones AP AP50 AP75 Baseline +DenoiseRep Baseline +DenoiseRep Baseline +DenoiseRep Mask-RCNN SwinV2-T 42.8% 44.3% 65.1% 67.1% 47.0% 48.6% SwinV2-S 48.2% 49.0% 69.9% 70.9% 52.8% 53.8% ResNet-50 42.6% 43.2% 63.7% 65.0% 46.4% 46.8% Faster-RCNN ResNet-50 37.4% 38.3% 58.1% 58.8% 40.4% 41.0% ATSS ResNet-50 39.4% 39.9% 57.6% 58.2% 42.8% 43.2% YOLO DarkNet-53 27.9% 28.4% 49.2% 50.3% 28.3% 27.8% DETR ResNet-50 39.9% 40.8% 60.4% 59.9% 41.7% 42.9% CenterNet ResNet-50 40.2% 40.6% 58.3% 59.1% 43.9% 44.0%
As shown in Table 8, we compare multiple classic backbone networks across different methods. After adding DenoiseRep, the accuracy index shows improvement without the need for additional model parameters. This indicates that our method enhances the representation learning capability of the model and extracts more effective features through denoising, all while maintaining the same time costs. In addition, DenoiseRep well to generalized to image detection tasks.
Appendix E Experiment on Image Segmentation Task
In this section, our objective is to assess the generalization capability of DenoiseRep in image segmentation tasks. The image segmentation task aims to divide an image into multiple regions in order to identify and understand objects or areas within them. The main challenges it faces include: complex and varied backgrounds that can easily interfere with segmentation results, objects obstruct each other, making segmentation difficult, objects have diverse shapes and may undergo deformations, etc. We conduct experiments on the ADE20K [zhou2017scene] dataset using the current mainstream image segmentation models. ADE20K is a widely used scene segmentation dataset, mainly used for image segmentation tasks. It contains approximately 20000 images and over 150 different object and region categories. We choose mIoU and B-IOU as evaluation metrics to comprehensively evaluate the performance of image segmentation models. MIoU is the average of IoUs for all categories, which can effectively reflect the segmentation ability of the model on different categories. It provides a quantitative model for the accuracy of handling complex scenes by calculating the degree of overlap between the predicted area and the real area. A higher mIoU value means that the model can better identify and segment the target object. B-IoU focuses on evaluating the accuracy of segmentation boundaries and is particularly suitable for object edge segmentation tasks. It provides sensitivity to boundary details by measuring the degree of overlap between predicted boundaries and real boundaries.
Table 9: The effectiveness of our method in image segmentation tasks was validated on ADE20K.
Methods Backbones aAcc B-IoU mIoU Baseline +DenoiseRep Baseline +DenoiseRep Baseline +DenoiseRep FCN [long2015fully] ResNet-50 0.774 0.779 0.287 0.299 0.359 0.365 FCN ResNet-101 0.793 0.796 0.306 0.316 0.396 0.404 SegFormer [xie2021segformer] mit_b0 0.782 0.788 0.292 0.297 0.374 0.381 SegFormer mit_b1 0.812 0.816 0.341 0.348 0.422 0.425
As shown in Table 9, we compare two classical backbone networks with different methods. After adding the DenoiseRep, both IoU metrics improve without adding model parameters. Practice proves that our method can improve the representation learning ability of the model and obtain more effective features through denoising without increasing additional time costs. In addition, DenoiseRep well to generalized to image segmentation tasks.
Appendix F Limitations